directions: Angle A and angle B are complementary angles. Find the measure of angle A and angle B.
measure angle A= 8x-7
measure angle B= x-11
i tried to place them equal to 90 degrees and i tried to put them equal to each other i don't think i got the correct answer with either way i tried.

Answer by richard1234(7193)

It means that their sum is equal to 90 degrees.

(8x - 7) + (x - 11) = 90

9x - 18 = 90

x = 108/9 = 12, so angle A = 89 and angle B = 1 (degree).